Under constant tension and constant mass per unit length, the note produced by a plucked string is 500 Hz when the length of the string is 0.90 m. At what length is the frequency 150 Hz?
1
Expert's answer
2020-11-05T10:33:03-0500
"f=\frac{1}{L}\sqrt{\frac{T}{m}}\\F_1L_1 = F_2L_2\\(500)0.9=150L_2\\L_2=3\ m"